OpenStudy (anonymous):

What is 3(x+1)=-2(x-1)+6. ?

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

give me a minute! :D just working this out

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

3x+3=-2x+2+6 3x+3=-2x+8 5x=5 x=1

OpenStudy (anonymous):

x=1(verified) :P

OpenStudy (anonymous):

What's the answer to -2(x+3)=-2(x-1)6

OpenStudy (anonymous):

2, 3 , all real numbers , or no solution

OpenStudy (anonymous):

(Help)

OpenStudy (anonymous):

?